# Tree-Wasserstein Distance for High Dimensional Data with a Latent Feature Hierarchy

### v2

@@ 1 Introduction
High-dimensional datasets with a latent feature hierarchy [Fefferman et al., 2016] are prevalent across various applications, involving data, e.g., genomic sequences [Tanay and Regev, 2017], text and images [Krishna et al., 2017], and movie user ratings [Bennett et al., 2007]. In this setting, the assumption is that there is an unknown underlying hierarchical structure to the data features, which is not directly observable. For example, in word-document data, where samples are documents and features are words, the hierarchy underlying features (words) is often assumed because words are naturally organized as nodes in a tree, each connected to others through various hierarchical linguistic relationships [Borge-Holthoefer and Arenas, 2010]. A key component of meaningful analyses of such data is finding distances between data samples that take into account the underlying hierarchical structure of features. As opposed to relying on generic metrics, e.g., the Euclidean or correlation distances, such meaningful distances serve as a fundamental building block in a broad range of downstream tasks, e.g., classification and clustering, often leading to significant performance improvement.
@@
Recently, hyperbolic geometry [Ratcliffe et al., 1994] has gained prominence in hierarchical representation learning [Chamberlain et al., 2017, Nickel and Kiela, 2017] because the lengths of geodesic paths in hyperbolic spaces grow exponentially with the radius [Sarkar, 2011], a property that naturally mirrors the exponential growth of the number of nodes in hierarchical structures as the depth increases. Methods using hyperbolic geometry typically focus on finding a hyperbolic embedding of the samples, relying on a (partially) known graph, whose nodes represent the samples [Sala et al., 2018]. However, considering such a known hierarchical structure of the samples is fundamentally different than the problem we consider here, where we aim to find meaningful distances between data samples that incorporate the latent hierarchical structure of the features.
@@
In this paper, we introduce a new tree-Wasserstein distance (TWD) [Indyk and Thaper, 2003] for this purpose, where we model samples as distributions supported on a latent hierarchical structure. We propose a two-step approach. In the first step, we embed features into continuous hyperbolic spaces [Bowditch, 2007] utilizing diffusion geometry [Coifman and Lafon, 2006] to approximate the hidden hierarchical metric underlying features. In the second step, we construct a tree that represents the latent feature hierarchy in a bottom-up manner from the hyperbolic embedding. The tree construction is based on a novel notion of a continuous analog of the lowest common ancestor (LCA) in trees defined in hyperbolic spaces by associating shortest paths on trees with hyperbolic geodesic distances. Utilizing the decoded feature tree, we introduce a new TWD for high-dimensional data with a latent feature hierarchy, based on diffusion geometry and hyperbolic geometry. We remark that in Lin et al. [2023], integrating hyperbolic and diffusion geometries was shown to be more effective in recovering latent hierarchies than existing hyperbolic embedding methods [Nickel and Kiela, 2017] (see App. E.3.4). However, the work in Lin et al. [2023] considers implicitly latent hierarchies of the samples, whereas here, we explicitly reveal the latent hierarchical structure of the features through a tree, which is then utilized for finding a meaningful TWD between the samples (see App. F.4). Using a TWD entails two useful attributes. First, the Wasserstein distance [Monge, 1781, Kantorovich, 1942] enables us to integrate feature relations into a meaningful distance between samples. Second, we use a tree to represent the latent feature hierarchy, learn this tree based on the data, and then, naturally incorporate it in the proposed TWD. Notably, this use of a tree significantly differs from the conventional use of trees in TWD for speeding up the computation of Wasserstein distance based on the Euclidean ground metric.
@@
From a theoretical standpoint, we show that our TWD provably recovers the Wasserstein distance based on the ground metric induced by the latent feature hierarchy. From a practical standpoint, we show that the proposed TWD is computationally efficient and scalable, making it suitable for large datasets. We showcase the efficacy of our method on an illustrative synthetic example, word-document datasets, and single-cell RNA-sequencing datasets. Specifically, we show that our TWD leads to superior classification performance compared to existing TWD-based methods [Indyk and Thaper, 2003, Backurs et al., 2020, Le et al., 2019, Yamada et al., 2022] and two application-specific methods based on pre-trained models (WMD [Kusner et al., 2015] and GMD [Bellazzi et al., 2021]).
@@ 2 Related Work
Wasserstein Distance. The Wasserstein distance [Villani, 2009] is a powerful tool for transforming the relations between features into a distance between probability distributions (samples from the probability simplex). It can be computed by solving the optimal transport (OT) problem [Monge, 1781, Kantorovich, 1942] with linear programming [Peyré et al., 2019] in cubic complexity. Various approaches, such as adding entropic regularization [Cuturi, 2013], using wavelet functions [Shirdhonkar and Jacobs, 2008, Gavish et al., 2010], applying Kantorovich-Rubinstein dual [Leeb and Coifman, 2016, Arjovsky et al., 2017], and computing the average Wasserstein distance across one-dimensional projections [Rabin et al., 2012], have been developed to reduce the computational complexity.
Tree-Wasserstein Distance. Another popular alternative to reduce the computational complexity of the Wasserstein distance is the tree-Wasserstein distance (TWD) [Indyk and Thaper, 2003], which can be computed in linear time, enabling efficient comparisons of large datasets. Existing TWD methods [Leeb, 2018, Sato et al., 2020, Takezawa et al., 2021, Yamada et al., 2022] attempt to approximate the Wasserstein distance based on the Euclidean ground metric with a tree metric by constructing a tree using the standard metric in the Euclidean ambient space (see App. B for more details). Our approach has a different goal. We assume that the features lie in a latent hierarchical space, infer a tree that represents this space, and compute a TWD based on the inferred tree to define a meaningful distance between the samples, incorporating the latent feature hierarchy.
Hyperbolic Representation Learning. While effective in computing embeddings and distances for hierarchical data samples, current hyperbolic representation learning methods [Chamberlain et al., 2017, Nickel and Kiela, 2017, 2018] are usually given a hierarchical graph, e.g., a tree, and do not include decoding a tree from observational data. One exception is the work in Chami et al. [2020], which introduces a method for decoding a tree in a two-dimensional Poincaré disk for hierarchical clustering [Murtagh and Contreras, 2012], by minimizing the continuous Dasgupta’s cost [Dasgupta, 2016]. However, this approach does not approximate the TWD between samples with a latent feature hierarchy because it does not recover the tree underlying the features.
@@ 3 Preliminaries
Tree-Wasserstein Distance. Consider a tree $T=([m],E,\mathbf{A})$ rooted at node 1 w.l.o.g., where $E$ is the edge set and $\mathbf{A}\in\mathbb{R}^{m\times m}$ contains edge weights. The tree metric $d_{T}$ between two nodes is the length of the shortest path on $T$. Let $\Gamma_{T}(v)$ be the set of nodes in the subtree of $T$ rooted at $v\in[m]$. For any $u\in[m]$, there exists a unique node $v$, which is the parent of $u$. We denote by $\omega_{u}=d_{T}(u,v)$ the length of the edge between any node $u$ and its parent $v$ (see App. A.4 for additional background on trees). The tree-Wasserstein distance (TWD) [Indyk and Thaper, 2003, Evans and Matsen, 2012, Le et al., 2019] between two probability distributions $\bm{\mu}_{1},\bm{\mu}_{2}\in\mathbb{R}^{m}$ supported on the tree $T$ is defined by
@@
Diffusion Geometry. Consider a set of points $\{\mathbf{a}_{j}\in\mathbb{R}^{n}\}_{j=1}^{m}$ assumed to lie on a manifold embedded in $\mathbb{R}^{n}$. Let $\mathbf{Q}\in\mathbb{R}^{m\times m}$ be an affinity matrix, given by $\mathbf{Q}_{jj^{\prime}}=\exp\left(-d^{2}(j,j^{\prime})/\epsilon\right)$, where $d(j,j^{\prime})$ is some suitable distance between $\mathbf{a}_{j}$ and $\mathbf{a}_{j^{\prime}}$, and $\epsilon>0$ is a scale parameter, often chosen as the median pairwise distance among data points scaled by a constant factor [Ding and Wu, 2020] (see App. D for hyperparameter tuning). Following Coifman and Lafon [2006], we consider the density-normalized affinity matrix $\widehat{\mathbf{Q}}=\mathbf{D}^{-1}\mathbf{Q}\mathbf{D}^{-1}$ to address the effects of non-uniform data sampling, where $\mathbf{D}$ is diagonal with entries $\mathbf{D}_{jj}=\textstyle\sum_{j^{\prime}}\mathbf{Q}_{jj^{\prime}}$, and the diffusion operator $\mathbf{P}$, defined by
@@
Poincaré Half-Space Model of Hyperbolic Geometry. Hyperbolic geometry is a Riemannian geometry with constant negative curvature $-1$ [Beardon, 2012]. We utilize the $m$-dimensional Poincaré half-space model $\mathbb{H}^{m}$, defined by $\mathbb{H}^{m}=\{\mathbf{x}\in\mathbb{R}^{m}|\mathbf{x}(m)>0\}$ with the Riemannian metric tensor $ds^{2}=(d\mathbf{x}^{2}(1)+d\mathbf{x}^{2}(2)+\ldots+d\mathbf{x}^{2}(m))/\mathbf{x}^{2}(m)$. The Riemannian distance for any two hyperbolic point $\mathbf{x},\mathbf{y}\in\mathbb{H}^{m}$ is defined by $d_{\mathbb{H}^{m}}(\mathbf{x},\mathbf{y})=2\sinh^{-1}(\left\lVert\mathbf{x}-\mathbf{y}\right\rVert_{2}/(2\sqrt{\mathbf{x}(m)\mathbf{y}(m)}))$.
Rooted Binary Tree. A rooted and balanced binary tree $B$ with $m$ leaves is a tree whose root node has degree 2, $m$ leaves have degree 1, and $m-2$ internal nodes have degree 3. For any two leaves $j,j^{\prime}$, let $j\vee j^{\prime}$ and $\Gamma_{B}(j\vee j^{\prime})$ denote their lowest common ancestor (LCA) and the set of leaves in the subtree of $B$ rooted at $j\vee j^{\prime}$, respectively. Given any three leaves $j_{1},j_{2},j_{3}$, the LCA relation $[j_{2}\vee j_{3}\sim j_{1}]_{B}$ holds if $\Gamma_{B}(j_{2}\vee j_{3})\subset\Gamma_{B}(j_{1}\vee j_{2}\vee j_{3})$ [Wang and Wang, 2018], as shown in Fig. 1(a). For binary trees, only one of $[j_{1}\vee j_{2}\sim j_{3}]_{B}$, $[j_{1}\vee j_{3}\sim j_{2}]_{B}$, or $[j_{3}\vee j_{2}\sim j_{1}]_{B}$ can hold.
@@ 4.1 Problem Formulation
Consider a data matrix $\mathbf{X}\in\mathbb{R}^{n\times m}$, whose $n$ rows consist of samples and $m$ columns consist of features. Suppose the features $\{\mathbf{X}_{:,1},\ldots,\mathbf{X}_{:,m}\}\subseteq\mathcal{H}\subset\mathbb{R}^{n}$ lie on an underlying manifold $\mathcal{H}$, which is a complete and simply connected Riemannian manifold with negative curvature embedded in a high-dimensional ambient space $\mathbb{R}^{n}$ with geodesic distance $d_{\mathcal{H}}$. We view $(\mathcal{H},d_{\mathcal{H}})$ as a hidden hierarchical metric space and consider a weighted tree $T=([m],E,\mathbf{A},\mathbf{X}^{\top})$ as its discrete approximation in the following sense. The tree node $j\in[m]$ is associated with the $j$-th feature, $E$ is the edge set, and $\mathbf{A}\in\mathbb{R}^{m\times m}$ is the weight matrix of the tree edges, defined such that the tree distance $d_{T}(j,j^{\prime})$ between two nodes $j$ and $j^{\prime}$, i.e., the length of the shortest path on $T$, coincides with the geodesic distance $d_{\mathcal{H}}(j,j^{\prime})$ between the features $j$ and $j^{\prime}$. Therefore, we refer to the hidden tree distance $d_{T}$ as the ground truth distance. Note that the assumption of an underlying manifold embedded in an ambient space $\mathbb{R}^{n}$ is widely used in studies of data manifolds and manifold learning [Shnitzer et al., 2022, Katz et al., 2020, Tong et al., 2021, Huguet et al., 2022, Kapuśniak et al., 2024], where the manifold typically underlies samples and graphs are viewed as discretizations of the underlying Riemannian manifold. In contrast, we consider the manifold underlying features. Specifically, we focus on a specification of the manifold to a hierarchical space $\mathcal{H}$ of features, and accordingly, a specification of the graph to a tree $T$. Our aim is to construct a tree $B$ from the data $\mathbf{X}$, such that the TWD between samples using our learned tree $B$ is (bilipschitz) equivalent to the Wasserstein distance using the ground truth latent tree distance $d_{T}$. To allow this construction of $B$, we assume that local affinities between (node) features $\{\mathbf{X}_{:,j}\}_{j=1}^{m}$ that are close in $\mathbb{R}^{n}$ embody information on the hidden hierarchical structure (see Sec. 4.4 for a formal description of this assumption).
Following a large body of work [Villani, 2009, Cuturi, 2013, Courty et al., 2017, Solomon et al., 2015, Yurochkin et al., 2019, Alvarez-Melis and Jaakkola, 2018], we focus on samples $\mathbf{X}_{i,:}\in\mathbb{R}^{m}_{+}$ with positive entries that can be normalized into discrete histograms $\mathbf{x}_{i}=\mathbf{X}_{i,:}/\left\lVert\mathbf{X}_{i,:}\right\rVert_{1}$, which is natural when the distribution over features is more important than the total mass. However, we note that our method is also applicable to data normalized into a histogram using other techniques.

### v3

@@ 1 Introduction
High-dimensional datasets with a latent feature hierarchy (Fefferman et al., 2016) are prevalent across various applications, involving data, e.g., genomic sequences (Tanay & Regev, 2017), text and images (Krishna et al., 2017), and movie user ratings (Bennett et al., 2007). In this setting, the assumption is that there is an unknown underlying hierarchical structure to the data features, which is not directly observable. For example, in word-document data, where samples are documents and features are words, the hierarchy underlying features (words) is often assumed because words are naturally organized as nodes in a tree, each connected to others through various hierarchical linguistic relationships (Borge-Holthoefer & Arenas, 2010). A key component of meaningful analyses of such data is finding distances between data samples that take into account the underlying hierarchical structure of features. As opposed to relying on generic metrics, e.g., the Euclidean or correlation distances, such meaningful distances serve as a fundamental building block in a broad range of downstream tasks, e.g., classification and clustering, often leading to significant performance improvement.
@@
Recently, hyperbolic geometry (Ratcliffe et al., 1994) has gained prominence in hierarchical representation learning (Chamberlain et al., 2017; Nickel & Kiela, 2017) because the lengths of geodesic paths in hyperbolic spaces grow exponentially with the radius (Sarkar, 2011), a property that naturally mirrors the exponential growth of the number of nodes in hierarchical structures as the depth increases. Methods using hyperbolic geometry typically focus on finding a hyperbolic embedding of the samples, relying on a (partially) known graph, whose nodes represent the samples (Sala et al., 2018). However, considering such a known hierarchical structure of the samples is fundamentally different than the problem we consider here, where we aim to find meaningful distances between data samples that incorporate the latent hierarchical structure of the features.
@@
In this paper, we introduce a new tree-Wasserstein distance (TWD) (Indyk & Thaper, 2003) for this purpose, where we model samples as distributions supported on a latent hierarchical structure. We propose a two-step approach. In the first step, we embed features into continuous hyperbolic spaces (Bowditch, 2007) utilizing diffusion geometry (Coifman & Lafon, 2006) to approximate the hidden hierarchical metric underlying features. In the second step, we construct a tree that represents the latent feature hierarchy in a bottom-up manner from the hyperbolic embedding. The tree construction is based on a novel notion of a continuous analog of the lowest common ancestor (LCA) in trees defined in hyperbolic spaces by associating shortest paths on trees with hyperbolic geodesic distances. Utilizing the decoded feature tree, we introduce a new TWD for high-dimensional data with a latent feature hierarchy, based on diffusion geometry and hyperbolic geometry. We remark that in Lin et al. (2023), integrating hyperbolic and diffusion geometries was shown to be more effective in recovering latent hierarchies than existing hyperbolic embedding methods (Nickel & Kiela, 2017) (see App. E.3.4). However, the work in Lin et al. (2023) considers implicitly latent hierarchies of the samples, whereas here, we explicitly reveal the latent hierarchical structure of the features through a tree, which is then utilized for finding a meaningful TWD between the samples (see App. F.4). Using a TWD entails two useful attributes. First, the Wasserstein distance (Monge, 1781; Kantorovich, 1942) enables us to integrate feature relations into a meaningful distance between samples. Second, we use a tree to represent the latent feature hierarchy, learn this tree based on the data, and then, naturally incorporate it in the proposed TWD. Notably, this use of a tree significantly differs from the conventional use of trees in TWD for speeding up the computation of Wasserstein distance based on the Euclidean ground metric.
@@
From a theoretical standpoint, we show that our TWD provably recovers the Wasserstein distance based on the ground metric induced by the latent feature hierarchy. From a practical standpoint, we show that the proposed TWD is computationally efficient and scalable, making it suitable for large datasets. We showcase the efficacy of our method on an illustrative synthetic example, word-document datasets, and single-cell RNA-sequencing datasets. Specifically, we show that our TWD leads to superior classification performance compared to existing TWD-based methods (Indyk & Thaper, 2003; Backurs et al., 2020; Le et al., 2019; Yamada et al., 2022) and two application-specific methods based on pre-trained models (WMD (Kusner et al., 2015) and GMD (Bellazzi et al., 2021)).
@@ 2 Related Work
Wasserstein Distance. The Wasserstein distance (Villani, 2009) is a powerful tool for transforming the relations between features into a distance between probability distributions (samples from the probability simplex). It can be computed by solving the optimal transport (OT) problem (Monge, 1781; Kantorovich, 1942) with linear programming (Peyré et al., 2019) in cubic complexity. Various approaches, such as adding entropic regularization (Cuturi, 2013), using wavelet functions (Shirdhonkar & Jacobs, 2008; Gavish et al., 2010), applying Kantorovich-Rubinstein dual (Leeb & Coifman, 2016; Arjovsky et al., 2017), and computing the average Wasserstein distance across one-dimensional projections (Rabin et al., 2012), have been developed to reduce the computational complexity.
Tree-Wasserstein Distance. Another popular alternative to reduce the computational complexity of the Wasserstein distance is the tree-Wasserstein distance (TWD) (Indyk & Thaper, 2003), which can be computed in linear time, enabling efficient comparisons of large datasets. Existing TWD methods (Leeb, 2018; Sato et al., 2020; Takezawa et al., 2021; Yamada et al., 2022) attempt to approximate the Wasserstein distance based on the Euclidean ground metric with a tree metric by constructing a tree using the standard metric in the Euclidean ambient space (see App. B for more details). Our approach has a different goal. We assume that the features lie in a latent hierarchical space, infer a tree that represents this space, and compute a TWD based on the inferred tree to define a meaningful distance between the samples, incorporating the latent feature hierarchy.
Hyperbolic Representation Learning. While effective in computing embeddings and distances for hierarchical data samples, current hyperbolic representation learning methods (Chamberlain et al., 2017; Nickel & Kiela, 2017; 2018) are usually given a hierarchical graph, e.g., a tree, and do not include decoding a tree from observational data. One exception is the work in Chami et al. (2020), which introduces a method for decoding a tree in a two-dimensional Poincaré disk for hierarchical clustering (Murtagh & Contreras, 2012), by minimizing the continuous Dasgupta’s cost (Dasgupta, 2016). However, this approach does not approximate the TWD between samples with a latent feature hierarchy because it does not recover the tree underlying the features.
@@ 3 Preliminaries
Tree-Wasserstein Distance. Consider a tree $T=(V,E,\mathbf{A})$ with $m$ leaves rooted at node 1 w.l.o.g., where $V$ is the vertex set, $E$ is the edge set and $\mathbf{A}\in\mathbb{R}^{\left|V\right|\times\left|V\right|}$ contains edge weights. The tree metric $d_{T}$ between two nodes is the length of the shortest path on $T$. Let $\Gamma_{T}(v)$ be the set of nodes in the subtree of $T$ rooted at $v\in V$. For any $u\in V$, there exists a unique node $v$, which is the parent of $u$. We denote by $\omega_{u}=d_{T}(u,v)$ the length of the edge between any node $u$ and its parent $v$ (see App. A.4 for additional background on trees). The TWD (Indyk & Thaper, 2003; Evans & Matsen, 2012; Le et al., 2019) between two distributions $\bm{\mu}_{1},\bm{\mu}_{2}\in\mathbb{R}^{m}$ supported on the tree $T$ is defined by
@@
Diffusion Geometry. Consider a set of points $\{\mathbf{a}_{j}\in\mathbb{R}^{n}\}_{j=1}^{m}$ assumed to lie on a manifold embedded in $\mathbb{R}^{n}$. Let $\mathbf{Q}\in\mathbb{R}^{m\times m}$ be an affinity matrix, given by $\mathbf{Q}_{jj^{\prime}}=\exp\left(-d^{2}(j,j^{\prime})/\epsilon\right)$, where $d(j,j^{\prime})$ is some suitable distance between $\mathbf{a}_{j}$ and $\mathbf{a}_{j^{\prime}}$, and $\epsilon>0$ is a scale parameter, often chosen as the median pairwise distance among data points scaled by a constant factor (Ding & Wu, 2020) (see App. D for hyperparameter tuning). Following Coifman & Lafon (2006), we consider the density-normalized affinity matrix $\widehat{\mathbf{Q}}=\mathbf{D}^{-1}\mathbf{Q}\mathbf{D}^{-1}$ to address the effects of non-uniform data sampling, where $\mathbf{D}$ is diagonal with entries $\mathbf{D}_{jj}=\textstyle\sum_{j^{\prime}}\mathbf{Q}_{jj^{\prime}}$, and the diffusion operator $\mathbf{P}$, defined by
@@
Poincaré Half-Space Model of Hyperbolic Geometry. Hyperbolic geometry is a Riemannian geometry with constant negative curvature $-1$ (Beardon, 2012). We utilize the $m$-dimensional Poincaré half-space model $\mathbb{H}^{m}$, defined by $\mathbb{H}^{m}=\{\mathbf{x}\in\mathbb{R}^{m}|\mathbf{x}(m)>0\}$ with the Riemannian metric tensor $ds^{2}=(d\mathbf{x}^{2}(1)+d\mathbf{x}^{2}(2)+\ldots+d\mathbf{x}^{2}(m))/\mathbf{x}^{2}(m)$. The Riemannian distance for any two hyperbolic point $\mathbf{x},\mathbf{y}\in\mathbb{H}^{m}$ is defined by $d_{\mathbb{H}^{m}}(\mathbf{x},\mathbf{y})=2\sinh^{-1}(\left\lVert\mathbf{x}-\mathbf{y}\right\rVert_{2}/(2\sqrt{\mathbf{x}(m)\mathbf{y}(m)}))$.
Rooted Binary Tree. A rooted and balanced binary tree $B$ with $m$ leaves is a tree whose root node has degree 2, $m$ leaves have degree 1, and $m-2$ internal nodes have degree 3. For any two leaves $j,j^{\prime}$, let $j\vee j^{\prime}$ and $\Gamma_{B}(j\vee j^{\prime})$ denote their lowest common ancestor (LCA) and the set of leaves in the subtree of $B$ rooted at $j\vee j^{\prime}$, respectively. Given any three leaves $j_{1},j_{2},j_{3}$, the LCA relation $[j_{2}\vee j_{3}\sim j_{1}]_{B}$ holds if $\Gamma_{B}(j_{2}\vee j_{3})\subset\Gamma_{B}(j_{1}\vee j_{2}\vee j_{3})$ (Wang & Wang, 2018), as shown in Fig. 1(a). For binary trees, only one of $[j_{1}\vee j_{2}\sim j_{3}]_{B}$, $[j_{1}\vee j_{3}\sim j_{2}]_{B}$, or $[j_{3}\vee j_{2}\sim j_{1}]_{B}$ can hold.
@@ 4.1 Problem Formulation
Consider a data matrix $\mathbf{X}\in\mathbb{R}^{n\times m}$, whose $n$ rows consist of samples and $m$ columns consist of features. Suppose the features $\{\mathbf{X}_{:,1},\ldots,\mathbf{X}_{:,m}\}\subseteq\mathcal{H}\subset\mathbb{R}^{n}$ lie on an underlying manifold $\mathcal{H}$, which is a complete and simply connected Riemannian manifold with negative curvature embedded in a high-dimensional ambient space $\mathbb{R}^{n}$ with geodesic distance $d_{\mathcal{H}}$. We view $(\mathcal{H},d_{\mathcal{H}})$ as a hidden hierarchical metric space and consider a weighted tree $T=([m],E,\mathbf{A},\mathbf{X}^{\top})$ as its discrete approximation in the following sense. The tree node $j\in[m]$ is associated with the $j$-th feature, $E$ is the edge set, and $\mathbf{A}\in\mathbb{R}^{m\times m}$ is the weight matrix of the tree edges, defined such that the tree distance $d_{T}(j,j^{\prime})$ between two nodes $j$ and $j^{\prime}$, i.e., the length of the shortest path on $T$, coincides with the geodesic distance $d_{\mathcal{H}}(j,j^{\prime})$ between the features $j$ and $j^{\prime}$. Therefore, we refer to the hidden tree distance $d_{T}$ as the ground truth distance. Note that the assumption of an underlying manifold embedded in an ambient space $\mathbb{R}^{n}$ is widely used in studies of data manifolds and manifold learning (Shnitzer et al., 2022; Katz et al., 2020; Tong et al., 2021; Huguet et al., 2022; Kapuśniak et al., 2024), where the manifold typically underlies samples and graphs are viewed as discretizations of the underlying Riemannian manifold. In contrast, we consider the manifold underlying features. Specifically, we focus on a specification of the manifold to a hierarchical space $\mathcal{H}$ of features, and accordingly, a specification of the graph to a tree $T$. Our aim is to construct a tree $B$ from the data $\mathbf{X}$, such that the TWD between samples using our learned tree $B$ is (bilipschitz) equivalent to the Wasserstein distance using the ground truth latent tree distance $d_{T}$. To allow this construction of $B$, we assume that local affinities between (node) features $\{\mathbf{X}_{:,j}\}_{j=1}^{m}$ that are close in $\mathbb{R}^{n}$ embody information on the hidden hierarchical structure (see Sec. 4.4 for a formal description of this assumption).
Following a large body of work (Villani, 2009; Cuturi, 2013; Courty et al., 2017; Solomon et al., 2015; Yurochkin et al., 2019; Alvarez-Melis & Jaakkola, 2018), we focus on samples $\mathbf{X}_{i,:}\in\mathbb{R}^{m}_{+}$ with positive entries that can be normalized into discrete histograms $\mathbf{x}_{i}=\mathbf{X}_{i,:}/\|\mathbf{X}_{i,:}\|_{1}$, which is natural when the distribution over features is more important than the total mass. However, we note that our method is also applicable to data normalized into a histogram using other techniques.